OpenStudy (anonymous):

Please help! Which of the following equations is written in standard form? A. -2x=13 B. -3x+13y=25

OpenStudy (anonymous):

I think none of them..

OpenStudy (anonymous):

By convention x coefficient will be positive while writing standard form..

OpenStudy (anonymous):

That's what I was thinking... there are two other answers but I'm positive they're wrong.

OpenStudy (anonymous):

Can you show us that??

OpenStudy (anonymous):

1/2x+1/3y=1/5 --- this wouldn't work because fractions aren't integers 3x+5y=-15 --- this also wouldn't work because the gcf isn't 1

OpenStudy (anonymous):

but, the gcf isn't 1..

OpenStudy (anonymous):

Standard form of an equation is given by: \[\large \color{green}{ax + by = c}\]

OpenStudy (anonymous):

Why are you looking at GCF ??

OpenStudy (anonymous):

because one of the rules of standard form is the gcf has to be 1!

OpenStudy (anonymous):

Yes you are right... It will be second one...

OpenStudy (anonymous):

ok. thanks.

OpenStudy (anonymous):

Welcome dear..